OpenStudy (anonymous):

What is the slope of the line 5x - 2y = -10? a. negative five halves b. five halves c. two fifths d. negative two fifths e. none of the above

imqwerty (imqwerty):

ok if we have a line of the form- ax+by+c=0 then its slope = -(a/b) so we have the equation -> 5x-2y=-10 = 5x-2y+10=0 comparing we see that a=5 and b=-2 so the slope = -(5/-2) =5/2 =2.5